Substance Q initially weighs 120% of its remaining mass after its decay in 60 minutes. What is its half-life in hours?


1
Expert's answer
2021-04-28T02:46:23-0400

Time = 60min = 1hr


Initial concentration= 1.2


Final concentration= 1


For first order reaction :-


"K =\frac {2.303} {t} log\frac {[initial]} {[final]}"


=> K = 0.18 hr-


Half life = "\frac {0.691} {K} =\frac {0.691} {0.18} = 3.8 hours"